OpenStudy (anonymous):

Please help! I need to solve for b. b=1/(2a), and....

OpenStudy (anonymous):

\[a=\sqrt{(2+\sqrt{5})/2}\] what is b?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } *\frac{ 1 }{\sqrt{(2+\sqrt{5})/2} }\]